OpenStudy (anonymous):

What is the sum of the first 5 terms in the geometric sequence a1=3 and r=1/3

OpenStudy (anonymous):

Whats your guess?

OpenStudy (anonymous):

i dont have a clue

OpenStudy (anonymous):

www.mathway.com

OpenStudy (anonymous):

helps you

OpenStudy (anonymous):

tried that

OpenStudy (anonymous):

this isnt an equation its a sequence

OpenStudy (anonymous):

A=3 and r=1/3

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what the sum of first five terms

OpenStudy (anonymous):

sum = a(1-r^n)/(1-r) =6[1-(1/3)^5]/(1-1/3) =6*(242/243)/(2/3) = 242/27

OpenStudy (anonymous):

was i right?